OpenStudy (anonymous):

What is the value of the expression when p = –24 and q = 4? Please help me I don't wanna fail 7th grade please help! p/2q

jimthompson5910 (jim_thompson5910):

you need to provide the expression

OpenStudy (jdoe0001):

got equation? * shamelessly taken from "got milk" *

OpenStudy (anonymous):

makes no sense....you are given p = -24 and q = 4....there is no p or q in your expression

OpenStudy (anonymous):

woops p/2q

OpenStudy (anonymous):

p/(2q) or (p/2)q......which one is it ?

OpenStudy (anonymous):

p/(2q)

OpenStudy (anonymous):

p/(2q) = -24/((2(4)) = -24/(8) = -3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW :)